OpenStudy (anonymous):

Evaluate the expression: (2 points) v ⋅ w Given the vectors: r = <8, 1, -6>; v = <6, 7, -3>; w = <-7, 5, 2>

OpenStudy (anonymous):

@misty1212 last one i promise ;)

OpenStudy (misty1212):

guess you can ignore the r part right?

OpenStudy (anonymous):

lol ye

OpenStudy (misty1212):

same as before, but now with three products instead of two

OpenStudy (anonymous):

please give me the answer i get to go to sleep after this and this is the last question

OpenStudy (misty1212):

\[6\times (-7)+7\times 5+(-3)\times 2\]

OpenStudy (anonymous):

71?

OpenStudy (anonymous):

@misty1212